TABLE OF CONTENTS

Publications in this Series . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . v Organizing Committees and Sponsors . . . . . . . . . . . . . . . . . . . . . . . . . vii Marcel Grossmann Awards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ix Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . xiii An Excursus on Experimental Gravitation From Space and Relativistic Astrophysics in Honour of Francis Everitt’s 60’th Birthday . . . . . . . . . . . . . . . . . . . . . . . . . . . . xv

PART A PLENARY SESSIONS

Chairperson: Humitaka Sato Was Einstein 100 Percent Right? Thibault Damour . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Relativistic Fluids and Gravitational Collapse Demetrios Christodoulou . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

Chairperson: David Finkelstein Do We Understand Black Hole Entropy? Jacob Bekenstein . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Unitary Rules for Black Hole Evaporation Andy Strominger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Recent Mathematical Developments in Quantum General Relativity Abhay Ashtekar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Black Hole and Particle-Like Solutions of the Einstein-Yang-Mills Equations Joel Smoller, Arthur G. Wasserman . . . . . . . . . . . . . . . . . . . . . . . 88 Gravitational Topological Charge and the Gravibreather Vladimir Belinsky . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 Quantum Gravity, the Planck Lattice and the Standard Model Giuliano Preparata . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

Chairperson: Abraham Taub The Lunar Orbit as Probe of Relativistic Gravity Ken Nordtvedt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 Gravitoelectromagnetism: Just a Big Word? Robert T. Jantzen, Paolo Carini, Donato Bini . . . . . . . . . . . . . . . . . 133

Approach to Gravitational Wave Astronomy Via the Regge-Wheeler-Teukolsky Equation Misao Sasaki . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

Chaos, Regularity, Noise in Self-Gravitating Systems Henry E. Kandrup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

Fully Relativistic 3D Numerical Simulations of Coalescing Binary Neutron Stars Takashi Nakamura . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183

Chairperson: Ray Weiss

The Advantages and Difficulties of Space Experiments for Testing General Relativity G. Mac Keiser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207

Real Performance and Real Promise—Resonant Gravitational Antennas Bill Hamilton . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

Chairperson: Jim Bardeen

COBE and Other Observations of CBR Anisotropy George Smoot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241

Primordial Nucleosynthesis as a Probe of Early Universe Katsuhiko Sato, N. Terasawa . . . . . . . . . . . . . . . . . . . . . . . . . 253

Cosmic Rays From 10^{17}ev to Beyond 10^{19}ev: Evidence From the Fly’s Eye Experiment and Ground Arrays Pierre Sokolsky . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264

PARALLEL SESSIONS

Exact Solutions Chairperson: V. Belinsky

Exact (3+1)-Dimensional Gravitational Soliton Solutions of the Einstein Equation Chi Au, Li-Zhi Fang, F.T. To . . . . . . . . . . . . . . . . . . . . . . . . . . 289

Radiating Dyon In Einstein-Maxwell Theory A. Chamorro, K.S. Virbhadra . . . . . . . . . . . . . . . . . . . . . . . . . 291

Geodesic Incompleteness and the Kinked De Sitter Spacetime K.A. Dunn, T.A. Harriott, J.G. Williams . . . . . . . . . . . . . . . . . . . 294

Bondi-Sachs Metrics and Exact Solutions S.J. Fletcher, A.W.-C. Lun . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 A Radiating Generalization of the Kerr Solution Dietrich Kramer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299

Cylindrical Black Hole in General Relativity and Black Holes in Lower Dimensional Gravity Theories José P.S. Lemos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302

Quantum Cosmology for the Bianchi Type IX Models Alfredo Macías, Michael P. Ryan, Jr. . . . . . . . . . . . . . . . . . . . . . 304

Metric of an Axisymmetric Neutron Star Vladimir S. Manko, J. Martin, E. Ruiz . . . . . . . . . . . . . . . . . . . . . 306

Dilaton Solutions with Arbitrary Electromagnetic Field Tonatiuh Matos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308

Extended N-Soliton Electrovac Solution E. Ruiz, V.S. Manko, J. Martin . . . . . . . . . . . . . . . . . . . . . . . . . 310

A New Exact Solution of Einstein Field Equations Yongjiu Wang, Zhiming Tang, Hongjun Guo, Hongwei Yu, Jiliang Jing . . . . . 313

Multi-Dimensional Gravitational Theories Chairperson: M. Demianski

Kaluza-Klein Solitons: New Singularities and Origin Definitions Andrew P. Billyard, Paul S. Wesson . . . . . . . . . . . . . . . . . . . . . . 317

Nonstandard-Dimensional Gravitational Theories Marek Demianski . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319

A Systematic Approach to the Study of Dimensional Reductions of Self-Dual Yang-Mills Equations Mario C. Díaz, George A.J. Sparling . . . . . . . . . . . . . . . . . . . . . . 325

The Role of the Internal Metric in Generalized Kaluza-Klein Theories W. Drechsler, D. Hartley . . . . . . . . . . . . . . . . . . . . . . . . . . . 328

The Classical Tests in (4+1)-Gravity D. Kalligas, P. Wesson, C.W.F Everitt . . . . . . . . . . . . . . . . . . . . 330

Space-Time-Matter Bahram Mashhoon, Hongya Liu, Paul S. Wesson . . . . . . . . . . . . . . . . 333

Fermi-Walker Coordinates in 2+1 Dimensional Gravity Pietro Menotti, Domenico Seminara . . . . . . . . . . . . . . . . . . . . . . 336

Variation of Fundamental Constraints and Red-Shift in 6-D Cosmological Models Americo Peraza Alvarez . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 Locally Supersymmetric Action in a D-Dimensional Superspace P. Salgado . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341

Toward Chiral Supergravity and Unification Kazunari Shima . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 343

Extreme Regimes of Gravitational Theory Chairperson: L. Halpern

Deformations, Large and Small, of Relativistic Membranes Riccardo Capovilla, Jemal Guven . . . . . . . . . . . . . . . . . . . . . . . 347

Highly Relativistic, Self-Gravitating Perfect Fluids with Differential Rotation F.J. Chinea . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 349

Is a Gravitational Geon Compatible With General Relativity? F.I. Cooperstock, V. Faraoni, G.P. Perry . . . . . . . . . . . . . . . . . . . 351

Extragalactic Background Radiation and the Time Symmetry of the Universe David A. Craig . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355

Stability Test for Cauchy Horizons T.M. Helliwell, D.A. Konkowski . . . . . . . . . . . . . . . . . . . . . . . . 358

Stability Tests for Mild Singularities D.A. Konkowski, T.M. Helliwell . . . . . . . . . . . . . . . . . . . . . . . . 360

Atomic Transitions Induced by Time-Dependent Spacetime Geometries Fabrizio Pinto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362

Approximation of Static Skyrmions from Connections in Compactified Minkowski Space M. Rosenbaum, A.A. Minzoni . . . . . . . . . . . . . . . . . . . . . . . . . . 366

Effects of Topology on the Gravitational Wave Jiro Soda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 369

Mathematical Foundations Chairperson: R. Schoen

Geometry Gravitation and Matter in Extreme Limits Leopold Halpern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 373

A Blunder in Quantum Field Theory Masayoshi Mizouchi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377

A Tensorial Lax Pair Equation and Integrable Systems in Relativity and Classical Mechanics Kjell Rosquist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 Stability of Non-Abelian Black Holes and Catastrophe Theory Takashi Tachizawa, Kei-Ichi Maeda, Takashi Torii . . . . . . . . . . . . . . . 386

The Role of Weyl Geometry in the Causal Interpretation of Quantum Mechanics W.R. Wood, G. Papini . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389

Quantum Probabilities Related to Wave Function’s Behavior with Respect to Spacetime Domain Norifumi Yamada . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 392

Mascroscopic Gravity Roustam M. Zalaletdinov . . . . . . . . . . . . . . . . . . . . . . . . . . . 394

Mathematical Relativity Chairperson: T. Dray

Algebraic Computing in General Relativity T. Dray . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401

A Class of Hyperbolic Gauge Conditions C. Bona, J. Stela, J. Massó, E. Seidel . . . . . . . . . . . . . . . . . . . . . 410

Mass and Angular Momentum from Taub Numbers Edward N. Glass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413

Ricci Fall-Off in Static, Globally Hyperbolic, Geodesically Complete, Ricci-Positive Spacetimes Steven G. Harris, David Garfinkle . . . . . . . . . . . . . . . . . . . . . . 415

MapleTensor: A New System for Performing Indicial and Component Tensor Calculations by Computer M. Kavian, R.G. McLenaghan, K.O. Geddes . . . . . . . . . . . . . . . . . . 418

Similarity Reductions of the Field Equations for Stationary Axisymmetric Rigidly Rotating Perfect Fluids Ben Langton, Edward Fackerell . . . . . . . . . . . . . . . . . . . . . . . . 420

MathTensor: Doing Tensor Analysis and Differential Forms by Computer Leonard Parker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 422

Static Axisymmetric Approach to the Two Black Hole Head-on Collision Samuel Rocha de Oliveira . . . . . . . . . . . . . . . . . . . . . . . . . . . 426

Spinors, Jets, and the Einstein Equations C.G. Torre . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 431

Positive Energy Proof for Asymptotically Anti-de Sitter Spacetimes E. Woolgar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 434 Phase Space and Gravitation Chairperson: G. Bisnovatyi-Kogan

Vlasov Equation in Spacetime Tangent Bundle Howard E. Brandt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439

Chaos in Static Axisymmetric Spacetimes Kei-Ichi Maeda, Yasuhide Sota, Shingo Suzuki . . . . . . . . . . . . . . . . . 441

Global Properties of Energy Truncated Spheroidal Stellar Systems Maria Teresa Menna, Giuseppe Pucacco, Remo Ruffini . . . . . . . . . . . . 444

Stability of Dense Stellar Clusters Against Relativistic Collapse: Maxwellian Distribution Functions with Different Cutoff Parameters Marco Merafina, Gennady S. Bisnovatyi-Kogan, Remo Ruffini, Enrico Vesperini . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447

Landau Damping in Semi-Degenerate Gravitating Systems R. Ruffini, M. Capalbi, S. Filippi, J.G. Gao, L.A. Sanchez . . . . . . . . . . . 451

Motion of Spin-1/2 Particles in External Gravitational and Electromagnetic Fields Stamatis Vokos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454

Scale-Invariant Phase Space and the Conformal Group James T. Wheeler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 457

Spacetime Coarse Grainings in Relativistic Particle Motion John T. Whelan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460

Alternate Gravitational Theories Chairperson: Y. Fujii

Summary of the Session: Alternative Gravitational Theories Yasunori Fujii . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 465

A Bianchi-Type IX Cosmological Solution in Brans-Dicke Theory P.AG. Chauvet, J.L. Cervantes-Cota . . . . . . . . . . . . . . . . . . . . . . 467

General Relativity and Empirical Success R. Disalle, W.L. Harper, S.R. Valluri . . . . . . . . . . . . . . . . . . . . . 470

The Fokker de Sitter Gyroscopic Precession in (Ω, A_{μν})-Field Theory Wang Tak Fung, Alfred Xin Yu . . . . . . . . . . . . . . . . . . . . . . . . 472

Gravitation and Spin, Minimal Coupling and Beyond Richard T. Hammond . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474

How Does Generalization of Einstein Theory Enhance Growth of Density Perturbations Toshinari Hirai, Kei-Ichi Maeda . . . . . . . . . . . . . . . . . . . . . . . . 477 Observational Limits on Scalar-Tensor Theories D. Kalligas, K. Nordtvedt, R.V. Wagoner . . . . . . . . . . . . . . . . . . . 480

The Gravitational Radiation of Ω-Field Theory Yu Yiu Lam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484

Higgs Mechanism and the Structure of the Energy-Momentum Tensor in Einstein Gravity and Conformal Gravity Philip D. Manneheim, Demosthenes Kazanas . . . . . . . . . . . . . . . . . . 486

The Equivalence Principle, Mach’s Principle and a Possible Link Between Lorentz-Invariant Scalar-Field Theory of Gravity and General Relativity P. Mazilu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489

Gravity Coupled to Vector Coulomb Fields Shin-ichi Nakariki, Kazumi Fukuma, Tetsuo Fukui, Masayoshi Mizouchi, Teruya Ohtani . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491

Crossing Symmetry is Incompatible with General Relativity H. Pierre Noyes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493

Quantum Behavior in Asymmetric, Weyl-Like Cartan Geometries J.E. Rankin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496

Isotopic Quantization of Gravity and its Universal Isopoincaré Symmetry Ruggero Maria Santilli . . . . . . . . . . . . . . . . . . . . . . . . . . . . 500

Motion of a Spinning Test Particle in the Ω-Field Metric Judy Tam Hong . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506

Inflation in Ω-Field Theory Lee Wa Tsan, Alfred Xin Yu . . . . . . . . . . . . . . . . . . . . . . . . . . 509

Advances on a Unified Physics-Astrophysics Theory Based on a Standing-Wave Particle Model Rafael A. Vera . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511

Can Gravity Explain Galactic Mysteries? Vadim V. Zhytnikov . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514

Gravitomagnetism Chairperson: D. Theiss

Gravitoelectromagnetism: Further Applications Donato Bini, Paolo Carini, Robert T. Jantzen . . . . . . . . . . . . . . . . . 519

Gravitoelectromagnetism and Inertial Forces in Relativity Paolo Carini, Donato Bini, Robert T. Jantzen . . . . . . . . . . . . . . . . . 522

The Question of Linearized Gravity and Maxwell-Newtonian Approximation C.Y. Lo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 On The Electromagnetism-Gravitation Analogy Nikolai V. Mitskievich . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528

The Gravitational and Electrostatic Fields Far from an Isolated Einstein-Maxwell Source J.H. Ocariz, Hector Rago . . . . . . . . . . . . . . . . . . . . . . . . . . . 530

Thomas Precession in Spacetime Geometries Abraham A. Ungar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 532

Perturbations of Stars and Black Holes Chairperson: V. Ferrari

Nonradial Pulsations of Compact Objects in General Relativity: An Overview Valeria Ferrari . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 537

Aspects of Black-Hole Scattering Nils Anderson . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560

Transformations Between the Regge-Wheeler and Bardeen-Press Equations J.F.Q. Fernandes, A.W.-C Lun . . . . . . . . . . . . . . . . . . . . . . . . . 564

Discharge of the Electromagnetic Field of a Reissner-Nordstrom Black Hole Rhett Herman, William A. Hiscock . . . . . . . . . . . . . . . . . . . . . . 567

Classically Unstable Modes of an Isolated Black Hole Osamu Kaburaki . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570

Non-Radial Oscillations of a Slowly Rotating Relativistic Star Yasufumi Kojima . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572

Quasinormal Frequencies of Step Potentials Hans-Peter Nollert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574

Quasinormal Modes of Neutron Stars and Black Holes: Kin or Strangers? Hans-Peter Nollert, Markus Leins, Michael Söffel . . . . . . . . . . . . . . 577

Quasi-Normal Modes and Late Time Behaviour of Waves in Gravitational Systems W.-M. Suen, E.S.C. Ching, P.T. Leung, K. Young . . . . . . . . . . . . . . . . 580

Small Disturbances in Black Hole Magnetosphere Toshio Uchida . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584

Groups in General Relativity Chairperson: M.P. Ryan, Jr.

Spacetime Admitting a 3-Parameter Group of Homotheties Jaume Carot, Alicia M. Sintes . . . . . . . . . . . . . . . . . . . . . . . . . 589

SO(2,4) Gauge Symmetry and the Einstein-Cartan-Maxwell Equations D.M. Kerrick . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 592 Conformal Motions in Bianchi I Spacetimes D. B. Lortan, S.D. Maharaj . . . . . . . . . . . . . . . . . . . . . . . . . . 594

A Generalization of the Noether Theorem Nikolai V. Mitskievich, Samir A. Sidawi . . . . . . . . . . . . . . . . . . . . 597

A Generalization of Ashtekar and the MacDowell-Mansouri Proposals Octavio Obregon, J.A. Nieto, J. Socorro . . . . . . . . . . . . . . . . . . . 599

Bubbles with an 0(3)-Symmetric Scalar Field in Curved Spacetime Nobuyuki Sakai, Yoonbai Kim, Kei-Ichi Maeda . . . . . . . . . . . . . . . . . 601

Connections on Principal Bundles Over Space-Time with Structure Groups SL(2,C) and O(3,C) A.H. Taub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 604

Tensor Multinomials and Einstein-Maxwell Theory Chris Vuille . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 606

Numerical Relativity Chairperson: M. Choptuik

Numerical Instabilities Associated with Gauge Modes Miguel Alcubierre, Bernard F. Schutz . . . . . . . . . . . . . . . . . . . . 611

Modelling Moving Black Holes Gabrielle Allen, Miguel Alcubierre, Simon Farrar, Bernard F. Schutz, Lee Aston Wild . . . . . . . . . . . . . . . . . . . . . . 615

Rotating Black Hole Spacetimes Steven R. Brandt, Edward Seidel . . . . . . . . . . . . . . . . . . . . . . . 619

Masses of Spindle and Cylindrical Naked Singularities Takeshi Chiba, Ken-Ichi Nakao, Takashi Nakamura . . . . . . . . . . . . . . 622

Initial Data for Quasi-Circular Orbits of Black-Hole Binaries Gregory B. Cook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624

Finite Differencing on the Sphere Roberto Gómez, Philippos Papadopoulos, Jeffrey Winicour . . . . . . . . . . 626

Critical Trapping of a Spherical Scalar Field Sean A. Hayward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629

A 3-D Apparent Horizon Finder Joan Massó, Joseph Libson, Edward Seidel, Wai-Mo Suen . . . . . . . . . . . 631

Adaptative Mesh Refinement in Numerical Relativity Joan Massó, Edward Seidel, Paul Walker . . . . . . . . . . . . . . . . . . . 634 Singularity Avoidance in Numerical Black Hole Spacetimes Joan Massó, Peter Anninos, Greg Daues, Edward Seidel, Wai-Mo Suen …. 637

Numerical Integration of the Robinson-Trautman Equation D.A. Prager, A.W.-C. Lun . . . . . . . . . . . . . . . . . . . . . . . . . . . 640

3-D Numerical Relativity at NCSA Edward Seidel, Peter Anninos, Karen Camarda, Joan Massó, Wai-Mo Suen, Malcolm Tobias, John Towns . . . . . . . . . . . . . . . . . . 644

Event Horizons of Numerical Black Holes Wai-Mo Suen, Peter Anninos, David Bernstein, Steve Brandt, Joseph Libson, Joan Massó, Edward Seidel, Larry Smarr, Paul Walker …. 648

Interface Behaviour in an Adaptive Mesh for Hyperbolic Equations Lee Aston Wild, Miguel Alcubierre, Gabrielle Allen, Bernard F. Schutz . . 651

Do One Dimensional Numerical Models of Supernovae Have Anything to Do With the Real World? James R. Wilson, Douglas S. Miller, Ronald W. Mayle . . . . . . . . . . . . 654

Relativistic Binary Neutron Star Coalescence James R. Wilson, Grant J. Mathews . . . . . . . . . . . . . . . . . . . . . . 658

Large Scale Structure Theory Chairperson: W. Stoeger

Robertson-Walker Models with Bulk Viscosity in Brans-Dicke Theory Aroonkumar Beesham . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 665

Decoherence of Homogeneous and Isotropic metrics in the Presence of Massive Vector Fields Orfeu Bertolami, P.V. Moniz . . . . . . . . . . . . . . . . . . . . . . . . . 668

The Missing Mass Problem in a Multiconnected Hyperspace-Time Claude Gauthier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 672

Observational Constraints on Primordial Perturbation Spectra Stefan Gottlöber . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 674

Numerical and Classical Cosmology Chairperson: B. Berger

Developments in Numerical Cosmology Beverly K. Berger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 679

Numerical Simulation of Unpolarized Gowdy Spacetimes David Garfinkle, Beverly Berger, Carrie Swift, Vijaya Swamy, Vincent Moncrief, Boro Grubisic . . . . . . . . . . . . . . . . . . . . . . . 699 Colliding Gravitational Waves in Friedmann-Robertson-Walker Backgrounds J.B. Griffiths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 702

A Numerical Kinetic Theory Approach for Two-Fluid Cosmological Models H. Harleston, H. Quevedo, R.A. Sussman . . . . . . . . . . . . . . . . . . . 705

Continuous Time Dynamics and Iterative Maps of Expansion Normalized Variables for Mixmaster Cosmologies D.W. Hobill, T.D. Creighton . . . . . . . . . . . . . . . . . . . . . . . . . . 708

On Properties of Large-Scale Inhomogeneities of Gravitational and Scalar Fields in the Vicinity of Cosmological Singularity A.A. Kirillov . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711

Rotational Perturbations of a Friedmann Universe Christian Klein . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 713

Tolman Bondi Spacetimes from a Different Viewpoint D.R. Matravers, R. Maartens, N.P. Humphreys . . . . . . . . . . . . . . . . 716

A Quantum-Driven-Time (QDT) Quantization of the Taub Cosmology Warner A. Miller, Arkady Kheyfets . . . . . . . . . . . . . . . . . . . . . 718

On The General Behaviour of the Universe Near the Cosmological Singularity Giovanni Montani . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 722

The Contour of Integration in Quantum Cosmology and Two Simplicial Minisuperspace Models Peter A. Morse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 726

Chaotic Inflation and Baryogenesis in Supergravity H. Murayama, Hiroshi Suzuki, T. Yanagida, Jun’ichi Yokoyama . . . . . . . . . 729

Evolution of a Perturbed Anisotropic Universe Hyerim Noh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 733

Bianchi V Models in N=2, D=5 Supergravity Luis O. Pimentel, J. Socorro . . . . . . . . . . . . . . . . . . . . . . . . . . 735

Cylindrical Solutions in N=2, D=5 supergravity Luis O. Pimentel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 738

Weyl Equation in Gödel Type Universes Luis O. Pimentel, Abel Camacho, Alfredo Macías . . . . . . . . . . . . . . . 741

Inflation in a Planar Universe Hisa-Aki Shinkai, Kei-Ichi Maeda . . . . . . . . . . . . . . . . . . . . . . . . 744

Heat Transport in an Inhomogeneous Spherically Symmetric Universe Joseph Triginer, Diego Pavón . . . . . . . . . . . . . . . . . . . . . . . . . 747 Quantum Gravity Chairperson: R. Kallosh

Exact Duality in String Effective Action Eric Bergshoeff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 751

Canonical Formulation and Finiteness of N=1 Supergravity with Supermatter P.D. D’Eath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 753

A Noncritical String Approach to Black Holes, Time and Quantum Dynamics John Ellis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 755

Average Effective Potential in Quantum Gravity Roberto Floreanini . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 758

Pair Creating Black Holes Jerome P. Gauntlett . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 761

IWP Solutions in Dilation-Axion Gravity David Kastor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 764

Bosonic Physical States in N=1 Supergravity? Miguel E. Ortiz, S.M. Carroll, D.Z. Freedman, D.N. Page . . . . . . . . . . . 767

Interpretation of Exact Quantum Taub Model Solutions in Terms of Classical Instanton Solutions Michael P. Ryan, Jr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770

Quantum Effects in 2+1 Dimensional Black Holes Alan R. Steif . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 773

Knots and Quantum Gravity Chairperson: J. Baez

Knots and Quantum Gravity: Progress and Prospects John C. Baez . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 779

Moduli Space Cohomology and Wavefunctionals in 3D Quantum Gravity Roger Brooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798

Skein Relations From Homotopy Transformations in Chern-Simons Theory Bernd Brügmann . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 801

Geometries Induced by Surfaces: An Algorithm on the 3-Dimensional Simplex Lattice Junichi Iwasaki . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 803

Vassiliev Invariants and Quantum Gravity Louis H. Kauffman . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 805 Wilson Loop Coordinates for 2+1 Gravity R. Loll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 807

Topological Feynman Rules for QGD Hugo A. Morales-Tecotl, Carlo Rovelli . . . . . . . . . . . . . . . . . . . 811

Gravity and Quantum Theory Chairperson: S. Carlip

Gravity and Quantum Theory Steve Carlip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 817

New Diagrammatic Method for Quantum Field Theory in the Heisenberg Picture Mitsuo Abe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 823

Charge and Central Charge in Chern-Simons Field Theory Máximo Bañados . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 825

On the Hamiltonian Formulation of Higher Dimensional Chern-Simons Gravity Máximo Bañados, Luis J. Garay . . . . . . . . . . . . . . . . . . . . . . . . 828

Anomaly in 2D-Gravity Coupled to Matter Fields Daniel Cangemi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 831

Fundamental Constants and the Problem of Time Alberto Carlini, Jeff Greensite . . . . . . . . . . . . . . . . . . . . . . . . 834

The Statistical Mechanics of the (2+1)-Dimensional Black Hole Steve Carlip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 837

Solution Generating Methods in Stringy Gravity Dmitri Gal’tsov, Alberto García . . . . . . . . . . . . . . . . . . . . . . . 840

On Black Hole Solutions in Multidimensional Gravitational Models V. D. Ivashchuk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 842

Finite Temperature Effects on Hyperbolic Space-Time K. Kirsten, G. Cognola, L. Vanzo, S. Zerbini . . . . . . . . . . . . . . . . . . 844

Mass Generation in Self-Interacting Ω-Field Theories K. Kirsten, E. Elizalde, G. Cognola, L. Vanzo, S. Zerbini . . . . . . . . . . . 846

Black Holes as Quantum Membranes Michele Maggiore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 848

The Spectral Analysis Inner Product for Quantum Gravity Donald Marolf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 851

Quantum Fluctuations of Black Hole Geometry Kouji Nakamura, Shigelu Konno, Yoshimi Oshiro, Akira Tomimatsu . . . . . . 854 Effective Potential and Dynamical Symmetry Breaking in 2D Induced Gravity Yu.I. Shil’nov, E. Elizalde, S.D. Odintzov, A. Romero . . . . . . . . . . . . . 857

The Einstein-Ostrogradsky-Dirac Hamiltonian Jonathan Z. Simon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 859

A General Supergravity Formalism for a Naturally Flat Inflation Potential Ewan D. Steward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 862

A Quadratic Spinor Lagrangian for General Relativity Roh-Suan Tung, James M. Nester . . . . . . . . . . . . . . . . . . . . . . . 865

String Theory in General Relativity Chairperson: L. Dixon

On a Possible Solution for the Polonyi Problem in String Cosmology M.C. Bento, O. Bertolami . . . . . . . . . . . . . . . . . . . . . . . . . . . 869

ALE Instantons in String Theory Massimo Bianchi, F. Fucito, G.C. Rossi, M. Martelli . . . . . . . . . . . . . . 873

Cherenkov Radiation From Superconducting Cosmic Strings Dmitri Gal’tsov, Yuri Grats . . . . . . . . . . . . . . . . . . . . . . . . . . 876

Stringy Black Holes from Sigma-Models Dmitri Gal’tsov, Alberto García, Oleg Kechkin . . . . . . . . . . . . . . . . 878

Black Holes and Sphalerons in Low Energy Effective String Theory George Lavrelashvili . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 880

Exact Four-Dimensional Dyonic Black Holes and Bertotti-Robinson Spacetimes in String Theory David A. Lowe, Andrew Strominger . . . . . . . . . . . . . . . . . . . . . . 883

Duality and Supersymmetry Tomás Ortín . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 889

Black Hole Entropy in Superstring Theory John Uglum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 891

PART B

Quantum Fields in General Relativity Chairperson: L. Parker

Effects of Quantum Fields on the Spacetime Geometries, Temperatures and Entropies of Static Black Holes Paul R. Anderson, Hooman Bahrani, William A. Hiscock, Janet Whitesell, James W. York, Jr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 897 On the Existence of Black Hole Evaporation V. Belinsky . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 900

Renormalization Group and Coarse-Graining in the Einstein Universe Alfio Bonanno . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 903

Radiation from a 2-D Evaporating Black Hole Sukanta Bose, Leonard Parker . . . . . . . . . . . . . . . . . . . . . . . . 908

Black Holes Can Change Mass When Nucleating Vacuum Phase Transitions Thomas J. Brueckner, William A. Hiscock . . . . . . . . . . . . . . . . . . . 915

Spinors Without Spinors Tevian Dray, Corinne A. Manogue, Jörg Schray . . . . . . . . . . . . . . . . 918

Averaged Energy Conditions and Quantum Inequality Restrictions on Negative Energy L.H. Ford, T.A. Roman . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 921

Relic Gravitons, Lightcone Fluctuations, and Ultraviolet Divergences L.H. Ford . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 924

Extreme Black Holes: A Third Venue for Quantum Gravity? William A. Hiscock, Daniel J. Loranz, Paul R. Anderson . . . . . . . . . . . 928

Graphical Representations of Invariants and Application to Quantum Gravity Shoichi Ichinose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 931

Two Regularization Methods of Energy-Momentum Tensor in Curved Space-Time Sang Pyo Kim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 934

Entropy Increase for Black Holes in Higher Curvature Gravity Robert Myers, Ted Jacobson, Gungwon Kang . . . . . . . . . . . . . . . . . 937

Stringy Twists of the Taub-Nut Metric Robert Myers, Clifford V. Johnson . . . . . . . . . . . . . . . . . . . . . . 940

Black Holes, the Wheeler-De Witt Equation and the Semi-Classical Approximation Miguel E. Ortiz . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 943

Chronology Protection and Quantized Field in Generalized Misner Space Tsunefumi Tanaka, William A. Hiscock . . . . . . . . . . . . . . . . . . . . 947

Topologically Induced Chaos in the Universe Roman Tomaschitz . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 950

Semiclassical Black Hole in Thermal Equilibrium with a Nonconformal Scalar Field Janet Whitesell, William A. Hiscock, Paul R. Anderson, James W. York, Jr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 952 Quantum Radiation and Accelerated Frames Chairperson: U. Gerlach

Paired Accelerated Frames Ulrich H. Gerlach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 957

The Equivalence Principle and The Boulware State in 1+1 Dimensions Warren G. Anderson . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 977

Quantum Cosmology Chairperson: D. Page

Probabilities Don’t Matter Don N. Page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 983

Inflation and Initial Singularities Arvind Borde . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1003

Physical States for Super Quantum Cosmology Riccardo Capovilla, Jemal Guven, Octavio Obregon . . . . . . . . . . . . . 1005

Quantization of Bianchi Models in N = 1 Supergravity with a Cosmological Constant A.D.Y. Cheng, P.D. D’Eath, P.R.L.V. Moniz . . . . . . . . . . . . . . . . . . 1007

Finite N = 1 Supergravity P.D. D’Eath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1010

Quantum Fluctuations of Planck Mass As Mutation Mechanism in a Theory of Evolution of the Universe Juan García-Bellido . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1012

Thermodynamic Stability of a Multi-Bubble Cosmological Model Gerald Horwitz, Oleg Fonarev . . . . . . . . . . . . . . . . . . . . . . . . 1014

Exact Solutions in Classical and Quantum Multidimensional Cosmology and Chaotic Behaviour Near the Singularity V.D. Ivashchuk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1018

Quantization of Bianchi Class A Models in Supergravity and the Probability Density Function of the Universe Hugh Luckock, Chris Oliwa . . . . . . . . . . . . . . . . . . . . . . . . . . 1020

Creation of a Child Universe by a False-Vacuum Bubble with Out-Going Negative Energy Radiation Takashi Mishima, Hiromi Suzuki, Noriaki Yoshino . . . . . . . . . . . . . . . 1023

Decoherence in Quantum Cosmology Masa-aki Sakagami, Takashi Okamura, Hiroto Kubotani . . . . . . . . . . . . 1026 Quantum State During and After the False Vacuum Decay Misao Sasaki, Takahiro Tanaka, Kazuhiro Yamamoto, Jun’ichi Yokoyama … 1028

Conformally Invariant Cosmology Yuri V. Shtanov . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1030

Quantum State After O(4)-Symmetric Bubble Nucleation with Gravitational Effects Takahiro Tanaka, Misao Sasaki . . . . . . . . . . . . . . . . . . . . . . . . 1032

Neutron Stars Chairperson: S. Tsuruta

Neutron Star Session Report: Some Selected Problems on Neutron Stars Sachiko Tsuruta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1037

Low-Frequency Modes of Pulsation of Relativistic Accretion Disks James R. Ipser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1048

The Effects of Superfluid Hydrodynamics on the Stability of Rotating Neutrons Stars Lee Lindblom, Gregory Mendell . . . . . . . . . . . . . . . . . . . . . . . . 1051

Self-Trapping of Superfluid Vortices and the Origin of Pulsar Glitches Yuko Mochizuki, Takeo Izuyama . . . . . . . . . . . . . . . . . . . . . . . . 1053

Dynamical Coherence in Neutron Stars G. Preparata, E. Del Giudice, R. Mele, C. Gualdi, G. Mangano, G. Miele … 1055

Neutron Stars with Strong Magnetic Fields Letao Qin, Sachiko Tsuruta . . . . . . . . . . . . . . . . . . . . . . . . . . 1064

Formation of Bosonic Compact Objects Edward Seidel, Wai-Mo Suen . . . . . . . . . . . . . . . . . . . . . . . . . . 1067

An Estimate of the Effect of Intermediate-Range Forces on the Structure of Neutron Stars Chris Vuille . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1070

Generation and Detection of Gravitational Waves Chairperson: J. Ipser

Modulation of Gravitational Waveforms from Merging Binaries Caused by Spin-Induced Orbital Precession Theocharis A. Apostolatos . . . . . . . . . . . . . . . . . . . . . . . . . . 1075

Compton Stochastic Gravitational Wave Background Katherine Compton, David Nicholson, Bernard F. Schutz . . . . . . . . . . . 1078 Efficient Gravitational Wave Chirp Detection and Estimation via Time-Frequency Analysis and Edge Detection M. Feo, V. Pierro, I.M. Pinto, M. Ricciardi . . . . . . . . . . . . . . . . . . 1086

Prospects for Deducing the Equation of State of Nuclear Matter from Gravitational-Wave Measurements of Neutron-Star Binary Coalescence Daniel Kennefick, Dustin Laurence, Kip S. Thorne . . . . . . . . . . . . . . 1090

Preparation for the Signal Search in the VIRGO Data B. Mours, B. Caron, A. Dominjon, F. Marion, L. Massonnet, R. Morand, M. Yvert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1093

A Possible Method for Estimating the Distances to Sources of Gravitational Waves Peter R. Saulson, William J. Startin . . . . . . . . . . . . . . . . . . . . . 1096

Cosmic Background Radiation Chairperson: S. Torres

Hot and Cold Spots in the Cobe-DMR Maps Laura Cayon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1101

Infrared Cutoff in Initial Density Perturbations and Cosmic Temperature Fluctuations Li-Zhi Fang, Yi-Peng Jing . . . . . . . . . . . . . . . . . . . . . . . . . . . 1104

Using Cosmic Variance and the COBE DMR Results to Constrain the Inflation Effective Potential Hannu Kurki-Suonio, Grant J. Mathews . . . . . . . . . . . . . . . . . . . . 1106

A New Calculation of the Sachs-Wolfe Effect and a Proof That Almost Isotropy of the Background Radiation Implenies the Universe Is Almost FLRW William R. Stoeger, S.J., George F.R. Ellis . . . . . . . . . . . . . . . . . . 1108

The Imprint of Ω on the Cosmic Microwave Background Naoshi Sugiyama . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1111

Gravitational Lenses Chairperson: A. Petters

Mathematical Aspects of Gravitational Lensing Arlie O. Petters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1117

Gravitational Lenses and Plastic Simulators Ronald J. Adler, William C. Barber, Mark E. Redar . . . . . . . . . . . . . 1138

The Odd Image Theorem Daniel Henry Gottlieb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1141

Effects of the Metric of Realistic Model Universes Mark W. Jacobs, Robert V. Wagoner . . . . . . . . . . . . . . . . . . . . . 1144 Gravitational Lensing Effect of Wormhole Sung-Won Kim, Y.M. Cho . . . . . . . . . . . . . . . . . . . . . . . . . . . 1147

Huygens’ Principle for Weyl’s Neutrino Equation and Maxwell’s Equations in Petrov Type III Space-Times R.G. McLenaghan, F.D. Sasse . . . . . . . . . . . . . . . . . . . . . . . . . 1149

Lensing in (4+1) Gravity James M. Overduin, Paul H. Lim, Paul S. Wesson . . . . . . . . . . . . . . . 1151

Criteria for Multiple Imaging in Lorentzian Manifolds Volker Perlick . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1153

Geometry and Matter in (4+1) Gravity Paul S. Wesson . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1156

Astroparticle Physics Chairperson: J. Frieman

Creation of a Scalar Potential in 2D Dilaton Gravity Klaus Behrndt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1161

Lasing Axion Clusters Thomas W. Kephart, Thomas J. Weiler . . . . . . . . . . . . . . . . . . . . 1164

Evolution Equations for Scale-Dependent Effective Actions M. Reuter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1168

A Solvable Model for Anharmonic Evolution of Cosmic Axion Overdensities Karl Strobl, Thomas J. Weiler . . . . . . . . . . . . . . . . . . . . . . . . 1171

Fluid Cosmology with Decay and Production of Particles Winfried Zimdahl, Diego Pavón . . . . . . . . . . . . . . . . . . . . . . . . 1175

Hubble Space Telescope Results Chairperson: D. Macchetto

Supernovae as Distance Indicators for Cosmology David Branch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1181

High Energy Phenomena at the Nuclei of Galaxies William B. Sparks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1184

Millisecond Pulsars Chairperson: M. Tavani

Millisecond Pulsars: Open Issues and Future Perspectives Marco Tavani . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1189

Timing of an Array of Millisecond Pulsars Donald C. Backer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1199 High Precision Metrology from Pulsar PSR J1713+0747 R.S. Foster, F. Camilo, A. Wolszczan . . . . . . . . . . . . . . . . . . . . . 1209

Searches for Millisecond Pulsars A.G. Lyne . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1213

Companion Evaporation in VLMXBS and Binary Millisecond Pulsars Jacob Shaham . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1223

Binary Pulsars and Neutron Star Masses S.E. Thorsett, Z. Arzoumanian . . . . . . . . . . . . . . . . . . . . . . . . . 1231

Compact Objects, Theory and Formation Chairperson: J. Wilson

Inner Structure of a Spherical Black Hole S. Droz, A. Bonnano, W. Israel, S.M. Morsink . . . . . . . . . . . . . . . . . 1239

Ernst-Like Formalism and New Differentially Rotating Solutions for Perfect Fluids in General Relativity L.M. González-Romero . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1242

Quasilocal Energy and Rotating Black Holes Erik A. Martinez . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1245

A Critique of the Expanding Photosphere Method Marcos J. Montes, Robert V. Wagoner . . . . . . . . . . . . . . . . . . . . 1247

Neutrino Flux Through the Conversion From Two-Flavor to Three-Flavor Quark Matter in a Supernova Core Quihe Peng, Zigao Dai, Tan Lu . . . . . . . . . . . . . . . . . . . . . . . . . 1250

The Gravitational Entropy of Collapsed Objects Quihe Peng . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1253

Axisymmetric Neutrino Radiation and the Mechanism of Supernova Explosions Tetsuya Shimizu, Shoichi Yamada, Katsuhiko Sato . . . . . . . . . . . . . . . 1256

Accreting Compact Objects Chairperson: L. Stella / P. Michelson

The Search for Black Holes in X-Ray Binaries: An Update Luigi Stella, G.L. Israel, S. Mereghetti, D. Ricci . . . . . . . . . . . . . . . 1261

Accretion of Matter with Magnetic Field onto a Black Hole G.S. Bisnovatyi-Kogan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1282

Highly Variable Magnetohydrodynamic Accretion onto a Black Hole Kouichi Hirotani, Akira Tomimatsu . . . . . . . . . . . . . . . . . . . . . . 1285 Relativistic Hadrons and the Origin of Relativistic Outflows in Active Galactic Nuclei Demosthenes Kazanas, John Contopoulos . . . . . . . . . . . . . . . . . . . 1287

Post-Newtonian Oscillations and Gravitational Radiation of a Rotating Disk of Dusk Wilhelm Kley, Gerhard Schäfer . . . . . . . . . . . . . . . . . . . . . . . . 1290

Quasar Redshifts Revisited From the Schwarzschild-Like Model W. Pitter, Z. Morantes, A. Aguillon, R. Urdaneta . . . . . . . . . . . . . . 1294

Processes of Matter Flow Nearby the Massive Compact Object Marina Romanova, Richard Lovelace . . . . . . . . . . . . . . . . . . . . . 1296

Trans-Fast MHD Accretion onto Black Hole Masaaki Takahashi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1298

Relativistic Diskoseismology: A New Signature of Black Holes Robert Wagoner, Michael Nowak, Christopher Perez . . . . . . . . . . . . . 1301

Problems in the Theory of Old Neutron Stars Accreting the Interstellar Medium S. Zane, A. Treves, M. Colpi, R. Turolla, L. Zampieri . . . . . . . . . . . . . 1304

Galactic X and γ Ray Sources Chairperson: P. Laurent / R. Ramaty

Very Short γ Bursts and Primordial Black Hole Evaporation David B. Cline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1309

Hard X-Ray Observations of Black Holes and Neutron Stars Philippe Laurent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1314

MACHOs and Dark Matter Chairperson: C. Stubbs

Production and Detection of Dark Matter Candidates: The PVLAS Experiment D. Bakalov, G. Cantatore, G. Carugno, S. Carusotto, P. Favaron, F. Della Valle, U. Gastaldi, E. Iacopini, E. Milotti, R. Onofrio, R. Pengo, F. Perrone, G. Petrucci, E. Polacco, C. Rizzo, G. Ruoso, E. Zavattini, G. Zavattini . . . . . . . . . . . . . . . . . . . . . . . . . . . 1329

A Possible Scenario for a Baryonic Dark Halo Francesco De Paolis, Gabriele Ingrosso, Philippe Jetzer, Marco Roncadelli . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1331

Dark Matter in NGC 4472 Francesco De Paolis, Gabrielle Ingrosso . . . . . . . . . . . . . . . . . . . 1333

Gravitational Microlensing and the Mass of the Dark Compact Halo Objects Philippe Jetzer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1336 Results from EROS (Expérience de Recherche d’Objets Sombres) Marc Moniez . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1338

Gravitational Wave Interferometers Chairperson: N. Kawashima

Functions of Proto-Type Laser Interferometers when KM Class Antenna Construction Started Nobuki Kawashima . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1345

Frequency Stabilization of a Nd:YAG Laser Using an Independently Suspended Fabry-Perot Cavity Akito Araya, K. Tsubono, N. Mio, S. Telada . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1347

Sensitivity of the LIGO Interferometer to Mirror Misalignment and Method for Automatic Alignment Yaron Hefetz, Nergis Mavalvala . . . . . . . . . . . . . . . . . . . . . . . . 1349

GEO 600: A 600 m Laser Interferometric Gravitational Wave Antenna Jim Hough, G.P. Newton, N.A. Robertson, H. Ward, A.M. Campbell, J.E. Logan, D.I. Robertson, K.A. Strain, K. Danzmann, H. Lück, A. Rüdiger, R. Schilling, M. Schrempel, W. Winkler, J.R.J. Bennett, V. Kose, M. Kühne, Bernard F. Schutz, D. Nicholson, J. Shuttleworth, H. Welling, P. Aufmuth, R. Rinkleff, A. Tünnermann, B. Willke . . . . . . . 1352

On a Possibility of Detecting Gravitational Waves Using Berry’s Phase for Photons Nikolai V. Mitskievich, Alexander I. Nesterov . . . . . . . . . . . . . . . . 1364

Development of 100m DL Type Laser Interferometer Gravitational Wave Antenna (TENKO-100) Ei-ichi Mizuno, N. Kawashima, R. Takahashi, S. Miyoki, K. Suehiro, S. Kawamura . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1366

On Topology Change in (2+1)-Dimensional Gravity Alexander I. Nesterov, Vladimir N. Efremov . . . . . . . . . . . . . . . . . 1368

A Recombined Fabry-Perot Prototype M. Ohashi, M.-K. Fujimoto, T. Yamazaki, M. Fukushima, A. Araya, S. Telada . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1370

Dilaton Black Holes with Electric Charge Malik Rakhmanov . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1372

Frequency Doubling of Nd:YAG Laser Light for Use in Gravitational Wave Detectors S. Rowan, A.M. Campbell, J. Hough, G.P. Newton, M. Gray, J. Hong . . . . . 1375

Frequency Locking of a ND:YAG Laser Using the Laser Itself as Optical Phase Modulator Giuseppe Ruoso, Guido Zavattini . . . . . . . . . . . . . . . . . . . . . . . . 1379 Status of the LIGO 40-Meter Interferometer R.L. Savage, Jr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1381

Scattering of Optical Components for Gravitational Wave Detectors Roland Schilling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1383

Laser Beam Geometry Fluctuations and Techniques for Their Suppression in Laser Interferometric Gravitiational Wave Detectors Kenneth Skeldon, K.A. Strain, H. Ward, J. Hough . . . . . . . . . . . . . . . 1385

Shot Noise Limited Sensitivity of the LIGO 40 Meter Interferometer Robert E. Spero . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1388

Imperfect Spherical Mirrors in the 100 m Delay-Line Interferometer R. Takahashi, S. Miyoki, E. Mizuno, E. Suehiro, S. Kawamura, N. Kawashima . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1390

The Glasgow Prototype Laser Interferometric Gravitational Wave Detector

— A Progress Report H. Ward. E Morrison, D.I. Robertson, J. Hough, S. Killbourn, G.P. Newton, N.A. Robertson, K.A. Strain . . . . . . . . . . . . . . . . . . 1392

Gravitational Wave Interferometers Chairperson: A. Giazotto

A New Technique for Low-Frequency Horizontal Vibration Isolation in the Suspension of a Laser-Interferometric Gravity Wave Detector Mark A. Barton, Kazuaki Kuroda . . . . . . . . . . . . . . . . . . . . . . . 1399

Concepts for Extending the Ultimate Sensitivity of Interferometric Gravitational Wave Detectors Using Non-transmissive Optics with Diffractive or Holographic Coupling R.W.P. Drever . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1401

Brownian Motion of a Torsion Pendulum Gabriela I. Gonzáles, Peter R. Saulson . . . . . . . . . . . . . . . . . . . . 1407

Towards the Development of Monolithic Fused Silica Suspensions for Laser Interferometric Gravitational Wave Detectors J.E. Logan, J. Hough, N.A. Robertson, K. Danzmann, R. Hutchins . . . . . . . 1410

Reduction of Outgassing Rate of Vacuum Ducts for Large-Scale Interferometric Gravitational-Wave Antennas: Application of Pre-Baking and Tin Coating N. Matuda, Y. Saito, Y. Ogawa, K. Akaishi, G. Horikoshi . . . . . . . . . . . 1413

AIGO: A Southern Hemisphere Second Generation Gravitational Wave Observatory D.E. McClelland, D.G. Blair, R.J. Sandman . . . . . . . . . . . . . . . . . . 1415

Toward the Development of a Second Generation Long Baseline Laser Interferometer for Gravitational Wave Detection D.E. McClelland, H.-A. Bachor, M.B. Gray, C.C. Harb, A.J. Stevenson …. 1418 Active/Passive Vibration Isolation at 1Hz and Above Part I: Single Stage, Six Degrees of Freedom D.B. Newell, S.J. Richman, R.T. Stebbins, P.G. Nelson, J.E. Mason, P.L. Bender, J.E. Faller . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1421

Operation of the Explorer Detector of the Rome Group Giovanni Vittorio Pallottino . . . . . . . . . . . . . . . . . . . . . . . . . 1424

Active/Passive Vibration Isolation at 1 Hz and Above Part II: Main Stages S.J. Richman, D.B. Newell, R.T. Stebbins, P.L. Bender, J.E. Faller . . . . . . 1426

Seismic Isolation in the LIGO 40 Meter Interferometer Robert E. Spero, L.A. Sievers . . . . . . . . . . . . . . . . . . . . . . . . . 1429

300-m Laser Interferometer Gravitational Wave Detector in Japan Kimio Tsubono . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1432

Measurement of Optical Path Fluctuations due to Residual Gas in the LIGO 40 Meter Interferometer M.E. Zucker, Stanley E. Whitcomb . . . . . . . . . . . . . . . . . . . . . . 1434

Gravitational Wave Bar Detectors Chairperson: D. Blair

Resonant Bar Gravitational Wave Detectors David G. Blair . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1439

A Capacitive B.A.E. Transducer Coupled to the ALTAIR Gravitational Wave Antenna: Results and Perspectives Paolo Bonifazi, Massimo Visco . . . . . . . . . . . . . . . . . . . . . . . . . 1441

Resonant Antenna for Monitoring Gravitational Wave Bursts in Our Galaxy Kazuaki Kuroda, Nobuyuki Kanda, Mark A. Barton, Jiro Arafune, Naoto Kondo, Norikatsu Mio, Kimio Tsubono . . . . . . . . . . . . . . . . . 1444

Low Noise Temperature Operation of the Niobium Gravitational Wave Antenna at the University of Western Australia Nick P. Linthorne, D.G. Blair, I.S. Heng, E.N. Ivanov, F. van Kann, M.E. Tobar, P.J. Turner . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1446

Optimal Reconstruction of the Input Signal in Resonant Gravitational Wave Detectors A. Ortolan, M. Biasotto, G. Maron, L. Taffarello, G. Vedovato, M. Cerdonio, G.A. Prodi, S. Vitale, J.P. Zendri . . . . . . . . . . . . . . . . 1459

An Interferometric Sapphire Transducer for Characterising Vibration Isolators H. Peng, L. Ju, D.G. Blair . . . . . . . . . . . . . . . . . . . . . . . . . . . 1461 Operation and Developments Concerning ALLEGRO-The Resonant Mass Antenna at Louisiana State University N. Solomonson, Z.K. Geng, W.O. Hamilton, W.W. Johnson, E. Mauceli, S. Merkowitz, A. Morse, B. Price . . . . . . . . . . . . . . . . . . . . . . . 1463

Signal-To-Noise Analysis for a Spherical Gravitational Wave Antenna Instrumented with Multiple Transducers Thomas R. Stevenson . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1466

Massive and Three-Mode Niobium-Sapphire Resonant-Bar Gravitational Wave Antennas Michael E. Tobar, D.G. Blair . . . . . . . . . . . . . . . . . . . . . . . . . 1469

Supernova 1987A Rome Maryland Gravitational Radiation Antenna Observations J. Weber . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1472

Gravitational Wave Bar Detectors Chairperson: M. Cerdonio

A 100 Ton 10mK Spherical Gravitational Wave Detector Giorgio Frossati . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1477

Solution of the Inverse Problem for a Spherical Gravitational Wave Antenna Nadja S. Magalhães, Warren Johnson, Carlos Frajuca, Odylio D. Aguiar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1481

Wideband Resonant-Lever Transducer for Massive Spherical Gravitational Wave Detectors Ho Jung Paik, Gregory M. Harry, Thomas Stevenson . . . . . . . . . . . . . 1483

The Ultracryogenic Gravitational Wave Antenna Nautilus G. Pizzella, P. Astone, M. Bassan, P. Bonifazi, P. Carelli, G. Castellano, E. Coccia, C. Cosmelli, V. Fafone, S. Frasca, A. Marini, G. Mazzitelli, I. Modena, P. Modestino, G.V. Pallottino, P. Rapagnani, F. Ricci, F. Ronga, M. Visco, L. Votano . . . . . . . . . . . . . 1486

Status of the Auriga G.W. Antenna G.A. Prodi, M. Cerdonio, L. Franceschini, G. Fontana, R. Mezzena, S. Paoli, S. Vitale, J.P. Zendri, M. Biasotto, M. Lollo, R. Macchietto, G. Maron, A. Ortolan, M. Strollo, G. Vedovato, M. Bonaldi, P. Falferi, E. Cavallini, P.L. Fortini, E. Montanari, L. Taffarello, A. Columbo, D. Pascoli, B. Tiveron . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1488

Multimode Optical Transducer for Massive Resonant Gravitational Wave Detectors Jean-Paul Richard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1490

Progress Towards a Two Mode Superconducting Transducer for Resonant Mass Gravitational Wave Detectors N. Solomonson, Z.K. Geng, W.O. Hamilton, W.W. Johnson . . . . . . . . . . . 1492 Behaviour of a DC Squid Tightly Couplened to a High-Q Resonant Transducer Thomas R. Stevenson, Hans J. Haucke . . . . . . . . . . . . . . . . . . . . . 1495

Lunar Laser Ranging and Space Geodesy Chairperson: R. Rheinhard / P.J. Shelus

Testing General Relativity with Lunar Laser Ranging John F. Chandler, Robert D. Reasenberg, Irwin I. Shapiro . . . . . . . . . . 1501

The Modified Hughes-Drever Experiment and the Strong Equivalence Principle Bi Ping Gong . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1505

A VLBI Measurement of the Solar Gravitational Deflection of Radio Waves D.E. Lebach, I.I. Shapiro, M.I. Ratner, J.L. Davis, B.E. Corey, T.A. Herring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1515

Determination of Relativistic Quantities by Analyzing Lunar Laser Ranging Data J. Müller, M. Schneider, M. Söffel, H. Ruder . . . . . . . . . . . . . . . . . 1517

A Mission Concept to Measure Second-Order Relativistic Effects, Solar Angular Momentum and Low-Frequency Gravitational Waves Wei-Tou Ni, An-Ming Wu, Jow-Tsong Shy . . . . . . . . . . . . . . . . . . . 1519

Laser Astrodynamics and Geodesy Wei-Tou Ni, Jow-Tsong Shy, Su-An Wu, Fu-Jeng Lin, Shie-Ei Wang, Sun-Kun King, An-Ming Wu, Shiao-Min Tseng, Ching-Yun Ren, Chang-Huei Wu, Chwen-Shell Ho, Ching-Ting Lee, Tze-Jang Chen . . . . . . 1522

General Relativity Parameters: The Relevance of Lunar Laser Data From the McDonald Laser Ranging Station Peter J. Shelus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1525

Frames of Reference in Relativistic Celestial Mechanics Slava G. Turyshev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1527

Relativity Parameters Determined from Lunar Laser Ranging James G. Williams, X.X. Newhall, Jean O. Dickey . . . . . . . . . . . . . . . 1529

Equivalence Principle and Inertial Frames Chairperson: J. Turneaure

The Relativity Mission Gyroscopes Saps Buchman, C.W. Francis Everitt, Bradford Parkinson, J.P. Turneaure, G.M. Keiser, Y. Xiao, D. Bardas, B. Muhlfelder, Dale Gill, Ping Zhou, P. Bayer, C. Gray, R. Brumley . . . . . . . . . . . . . 1533

Induced Composition-Dependence in BD Theory Yasunori Fujii . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1536 The Scientific Significance of Long-Lifetime High-Accuracy Version of the Relativity-Gyro Experiment with Altitude-Chance Capability Benjamin Lange . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1538

Status of the Cryogenic Inertial Reference System for the Gravity Probe B Mission J.A. Lipa, D.-H. Gwo, S. Wang, R.K. Kirschman . . . . . . . . . . . . . . . . 1542

A Status Report on the Relativity Mission B. Muhlfelder, C.W.F. Everitt, B.W. Parkinson, J.P. Turneaure, D. Bardas, S. Buchman, D.B. DeBra, H. Dougherty, D. Gill, G. Green, G.M. Gutt, J. Gwo, N.J. Kasdin, G.M. Keiser, T. Langenstein, J.A. Lipa, J.M. Lockhart, R.T. Parmely, M.A. Taber, R.A. Van Patten, Y.M. Xiao, S. Wang . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1545

Geodesy with Step Comparison of Various Mission Options J. Müller, H. Fröhlich, R. Rummel, N. Sneeuw . . . . . . . . . . . . . . . . . 1551

The Guide Star for the Gyroscope Relativity Mission Michael I. Ratner, D.E. Lebach, I.I. Shapiro, N. Bartel . . . . . . . . . . . . 1553

Gravitomagnetic Curvature Effects and the Motion of Local Inertial Frames Dietmar S. Theiss . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1555

Measurement of the Earth’s Gravitomagnetic Field Using Space Gradiometry Dietmar S. Theiss . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1558

Scaling Laws for Rotating Fluid Shapes Rodney H. Torii, J.E. McCuan, P.J. Neschleba, P.I. Lin, R. Finn . . . . . . . . 1561

Accelerometric Reference Sensor for Interferometer Space Antenna P. Touboul, A. Rüdiger . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1563

Servo-Controlled Accelerometers for Gravitational Experiments in Space P. Touboul, B. Foulon, A. Bernard . . . . . . . . . . . . . . . . . . . . . . 1565

Error Analysis for the QuickSTEP Experiment Paul W. Worden, Jr., Michael A. Swartwout . . . . . . . . . . . . . . . . . 1567

The QuickSTEP Instrument and Science Paul W. Worden, Jr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1569

Readout Schemes for the GP-B Experiment Y.M. Xiao, G.M. Keiser . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1574

Precision Gravitational Measurements Chairperson: P. Boynton

Progress Report on a Rotatable Torsion Balance for a New Test of the Equivalence Principle M.A. Beilby, N. Krishnan, R.D. Newman . . . . . . . . . . . . . . . . . . . . 1579 Equivalence Principle and Feeble Yukawa-like Forces Ramanath Cowsik, N. Krishnan, S.N. Tandon, C.S. Unnikrishnan . . . . . . . . 1583

The Velocity of Particles of Galactic Dark Matter Ramanath Cowsik, Charu Ratnam, P. Bhattacharjee . . . . . . . . . . . . . 1597

Precise Experimental Free Fall Tests of the Weak Principle of Equivalence H. Dittus, R. Greger, S.T. Lochmann, P. Mazilu . . . . . . . . . . . . . . . . 1599

Search for Fundamentally New Interactions with Ranges > 1cm J.H. Gundlach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1601

Cryogenic Torsion Balances — A Pilot Study N. Krishnan, M.K. Bantel, M.A. Beilby, A. Hunt, R.D. Newman . . . . . . . . 1604

Testing Gravity with Atom Beam Interferometry Claus Lämmerzahl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1607

Solar-Neutrino Experiments with Null Results J. Luo, X. Chen, J.G. Li, S.H. Fan, F.Y. Li . . . . . . . . . . . . . . . . . . . 1610

Detection of Non-Newtonian Forces M.W. Moore, P.E. Boynton . . . . . . . . . . . . . . . . . . . . . . . . . . 1612

Future Torsion Balances for Gravitation Experiments: What Are Their Limits? Riley D. Newman, M. Bantel, M. Beilby, N. Krishnan, E. Siragusa, P.E. Boynton, A. Goodson . . . . . . . . . . . . . . . . . . . . . . . . . . . 1619

Search For Anomalous Spin-Dependent Forces Using a DC Squid With a Paramagnetic Salt Wei-Tou Ni, Sheau-Shi Pan, Hsien-Chi Yeh, Li-Shing Hou, Juling Wan . . . . . 1625

Test of Quantum Electrodynamics and Search For Light Scalar/Pseudoscalar Particles Using Ultra-High Sensitive Interferometers Wei-Tou Ni, Sun-Kun King, Hung-Wen Chen, Jow-Tsong Shy, Norikatsu Mio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1628

Photon Fall in a Gravitational Field at LIGO James T. Wheeler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1631

Space Gravitational Wave Experiments Chairperson: R. Hellings

Pioneer 10 1993 Gravitational Wave Experiment J.D. Anderson, E.L. Lau, J.W. Armstrong . . . . . . . . . . . . . . . . . . . 1637

Search for Gravitational Wave By Doppler Tracking of Interplanetary Spacecraft Bruno Bertotti . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1638

LISA: Laser Interferometer Space Antenna for Gravitational Wave Measurements Karsten Danzmann . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1641 Thermal Stability Analysis for a Heliocentric Gravitational Radiation Detection Mission W. M. Folkner, P. McElroy, R. Miyake, P.L. Bender, R.T. Stebbins, W. Supper . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1655

Cassini Relativity Experiments Giacomo Giampieri . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1658

Optics for Laser Transponders D.I. Robertson, J. Hough, A. McLaren, H. Ward, K. Danzmann . . . . . . . . 1665

Algorithms for Unequal-Arm Michelson Interferometers Massimo Tinto, Giacomo Giampieri, Ronald W. Hellings, Peter Bender, James E. Faller . . . . . . . . . . . . . . . . . . . . . . . . 1668

One-Arm Interferometric Detectors of Gravitational Waves Massimo Tinto, Frank B. Estabrook . . . . . . . . . . . . . . . . . . . . . . 1671

PUBLIC TALKS

Copernican Humility, Change and the Creation of Purpose Yeval Ne’eman . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1677

The Physical Heritage of the Cold War Wolfgang K.H. Panofsky . . . . . . . . . . . . . . . . . . . . . . . . . . . 1717

TEST: Traction of Events in Space-Time Anna Imponente . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1738

List of Participants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1744

Author Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1760